Tim and Pam are 88 miles apart. At the same time, each starts traveling toward the other. Time travels at 30 mph. Pam travels at 14 mph. After how long do they meet?
How do you solve these types of problems?
thank you for your help

Answer by ankor@dixie-net.com(15624)

Tim and Pam are 88 miles apart. At the same time, each starts traveling toward the other.
Tim travels at 30 mph. Pam travels at 14 mph. After how long do they meet?
:
Actually you can do this one in your head. Since they are going toward each other,
their relative speed is 30 + 14 = 44 mph
How long to cover 88 mi at 44 mph? 2 hrs, right
;
If you want an equation, write a distance equation: dist = speed * time
Let t = time for them to meet
:
Tim's dist + Pam's dist = 88 mi
30t + 14t = 88
44t = 88
t = 88/44
t = 2 hrs